# Computable Thermodynamic Formalism

### v1

@@ 5.1. Jacobian and the transfer operator
Finally, we show that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. We argue this for $n$ by contradiction and consider a measure $\mu\in\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}$ with $\mu\neq\mu_{n}$. Then it is not hard to derive from (5.29) that $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ by item (iv) of the Assumptions. Thus, since $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$, we have $\mu(\{x_{0}\})>0$ for some $x_{0}\in C_{n}\smallsetminus Y_{n}$. Then we show that $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. By the assumptions in Theorem 5.9, $p_{n,i}\in U_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Hence, by $\mu\in\mathcal{K}_{n}$, it follows from (5.28) that $0\leqslant\mu(\{p_{n,i}\})\leqslant\mu(U_{n,i,k})\leqslant t_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Note that $\inf_{k\in\mathbb{N}}t_{n,i,k}=0$ for each $(n,i)\in H_{n}$ by (5.27). Then we have $\mu(\{p_{n,i}\})=0$ for each $(n,i)\in H_{n}$. Hence, by item (v) of the Assumptions, we obtain that $\mu(C_{n}\smallsetminus X_{n})=\mu(B_{n})=0$. Note that $\mu\in\mathcal{P}(X;C_{n})\cap\mathcal{M}(X,T_{n};X_{n})$ by $\mu\in\mathcal{K}_{n}$ and (5.28). Then by (1.1), we have $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n};X_{n})=\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. Hence, by Lemma 5.8, it follows from $\mu(\{x_{0}\})>0$ that $x_{0}$ is a periodic point of $T_{n}$ in $C_{n}\cap(X_{n}\smallsetminus Y_{n})$. Moreover, by $\mu\in\mathcal{P}(X;C_{n}\cap X_{n})$, the orbit of $x_{0}$ for $T_{n}$ is contained in $C_{n}\cap X_{n}$. Thus, $x_{0}\in B_{n}$ by the definition of $B_{n}$. This implies that $\mu(B_{n})\geqslant\mu(\{x_{0}\})>0$, which contradicts that $\mu(B_{n})=0$. Hence, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. This completes the proof of Claim 2.
@@
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions whose maps $T_{n}$ are uniformly contracting near all periodic orbits in $B_{n}$ for $T_{n}$. Note that for a rational map, all periodic orbits containing critical points are super-attracting. Hence, Theorem 5.10 can be applied to establish the computability of equilibrium states for rational maps and potentials satisfying certain hypotheses.
@@
    There exists a computable function $m\colon H\rightarrow\mathbb{N}$ and two sequences $\{r_{n,i}\}_{(n,i)\in H}$, $\{\lambda_{n,i}\}_{(n,i)\in H}$ of uniformly computable real numbers such that for each $(n,i)\in H$, $m(n,i)$ is the period of $p_{n,i}$ for $T_{n}$, and
    $\rho\bigl(p_{n,i},(T_{n})^{m(n,i)}(q)\bigr)\leqslant\lambda_{n,i}\cdot\rho(p_{n,i},q)\quad\text{for each }q\in B(p_{n,i},r_{n,i}).$
@@
We establish that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$ for each $(n,i)\in H$. We consider an arbitrary pair $(n,i)\in H$ and define $B(l)\coloneqq(T_{n})^{lm(n,i)}(B(p_{n,i},r_{n,i}))$ for each $l\in\mathbb{N}_{0}$.
By item (iv) of the Assumptions, it follows from $C_{n}\subseteq X_{n}$ that $p_{n,i}\in B_{n}\subseteq C_{n}\cap(X_{n}\smallsetminus Y_{n})=C_{n}\smallsetminus Y_{n}$. Note that $\mu_{n}\in\mathcal{P}(X;C_{n}\cap Y_{n})$ by item (iv) of the Assumptions. Then we have $\mu_{n}(\{p_{n,i}\})=0$. Moreover, by the definition of $\{B(l)\}_{l\in\mathbb{N}_{0}}$, it follows from $\mu_{n}\in\mathcal{E}_{0}(T_{n},\phi_{n})\subseteq\mathcal{M}(X,T_{n})$ that $\mu_{n}(B(0))\leqslant\mu_{n}(B(l))$ for each $l\in\mathbb{N}_{0}$. Since $\lambda_{n,i}\in(0,1)$, we have $B(l)\subseteq B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\subseteq B(0)$ for each $l\in\mathbb{N}_{0}$. Thus, we have $\mu_{n}(B(0))=\mu_{n}\bigl(B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=\mu_{n}(B(l))$, hence, $\mu_{n}\bigl(B(0)\smallsetminus B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=0$. Letting $l\rightarrow+\infty$, we obtain that $\mu_{n}(B(0)\smallsetminus\{p_{n,i}\})=0$. Combined with $\mu_{n}(\{p_{n,i}\})=0$, this proves that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$.
@@ 6. Computability of equilibrium states for expanding Thurston maps
In this section, we consider a class of non-uniformly expanding maps on the topological $2$-sphere known as expanding Thurston maps. There has been active research on both thermodynamic formalism for expanding Thurston maps (see e.g. [BM10, BM17, HP09, Li18, Li15, Li17, LS24b]) and algorithmic aspects of these maps (see e.g. [SY15, RSY20]).
We first review the definition of expanding Thurston maps, along with some key concepts and results. Then in Subsection 6.2, we focus on Misiurewicz–Thurston rational maps and apply Approach I (see Theorem 1.1) to establish Theorem 1.2. Finally, in Subsection 6.3, we study computable expanding Thurston maps whose critical points are computable. By Theorem 6.4, their measure-theoretic entropy functions may not be upper semi-continuous, which prevents us from using Approach I (see Theorem 1.1) to demonstrate the computability of their corresponding equilibrium states. Instead, we apply Approach II (see Theorem 1.3) to prove Theorem 1.4.
@@ 6.1. Expanding Thurston maps
$\lim_{n\to+\infty}\sup\bigl\{\operatorname{diam}_{d}X:X\text{ is a connected component of }f^{-n}\bigl(S^{2}\smallsetminus\mathcal{C}\bigr)\bigr\}=0.$ (6.1)
@@
The following result regarding the upper semi-continuity of the measure-theoretic entropy function is established in [LS24b, Theorem 1.1], extending [Li15, Corollary 1.3].
@@
Let $f\colon S^{2}\rightarrow S^{2}$ be an expanding Thurston map. Then the measure-theoretic entropy function of $f$ is upper semi-continuous if and only if $f$ has no periodic critical points.
@@ 6.2. Misiurewicz–Thurston rational maps
For all $n,\,m\in\mathbb{N}$, $\phi,\,u\in C(\widehat{\mathbb{C}},\sigma)$, a Misiurewicz–Thurston rational map $f=h_{1}/h_{2}$, and $x\in\widehat{\mathbb{C}}\smallsetminus\bigl\{f^{i}(\infty):1\leqslant i\leqslant m\bigr\}$, the algorithm outputs a rational $2^{-n}$-approximation of $\mathcal{L}^{m}_{\phi}(u)(x)$, given the following input data:
@@
Since there exists no periodic critical point of the Misiurewicz–Thurston rational map $f$, by Theorem 6.4, the measure-theoretic entropy function $\nu\mapsto h_{\nu}(f)$ is upper semi-continuous. By the hypothesis of Theorem 1.2, we have $\mathcal{E}(f,\phi_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$. Hence, to prove Theorem 1.2, by Theorem 1.1, it suffices to verify that the sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of functions $\phi_{n}\colon\widehat{\mathbb{C}}\rightarrow\mathbb{R}$ satisfies properties (i) and (ii) in Theorem 1.1 in the case where $T_{n}\coloneqq f$ for each $n\in\mathbb{N}$.
@@
By (6.5), $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\big\lvert}q_{k,t}\mathclose{\big\rvert}$ for each $k\in\mathbb{N}$. Hence, by Definition 3.12, the function $F\colon\mathbb{N}\rightarrow\mathbb{R}$ given by $F(k)\coloneqq\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\big\lvert}q_{k,t}\mathclose{\big\rvert}$ for each $k\in\mathbb{N}$ is a computable function satisfying $F(k)\geqslant\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}$ for each $k\in\mathbb{N}$. Thus, by (6.4), we have $\psi_{k}\in C^{0,\beta}\bigl(\widehat{\mathbb{C}},d_{v}\bigr)$ and $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{\beta,d_{v}}\leqslant C\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant CF(k)$ for each $k\in\mathbb{N}$. By Proposition 6.8, $\{P(f,\psi_{k})\}_{k\in\mathbb{N}}$ is a sequence of uniformly computable real numbers. This completes the proof of the claim.
@@ 6.3. Computable expanding Thurston maps
Before the proof of Theorem 1.4, we establish the following results.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Suppose that all balls in $(X,\rho)$ are connected. Assume that $I$ is a recursively enumerable set and that $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then there exists a recursively enumerable set $E\subseteq\mathbb{N}\times I$ and a sequence $\{V_{n,i}\}_{(n,i)\in E}$ of uniformly lower semi-computable open sets that are connected such that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$, where $E_{i}\coloneqq\{(n,i)\in E:n\in\mathbb{N}\}$ for each $i\in I$.
@@
Now we fix $i\in I$ and check that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$. Indeed, for each $(n,i)\in E_{i}$, since $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)$ is connected and $s_{n}\in U_{i}$, we have $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)\subseteq U_{i}$. By the definition of $\{V_{n,i}\}_{(n,i)\in E}$, this implies that $\bigcup_{(n,i)\in E}V_{n,i}\subseteq U_{i}$. Then we establish $U_{i}\subseteq\bigcup_{(n,i)\in E_{i}}V_{n,i}$. We consider an arbitrary $x_{0}\in U_{i}$, and show that $x_{0}\in V_{n_{0},i}$ for some $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$. Since $\{s_{n}:n\in\mathbb{N}\}$ is dense in $X$ and $\{s_{n}:(n,i)\in E_{i}\}=\{s_{n}:n\in\mathbb{N}\}\cap U_{i}$, there exists $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$ such that $\rho\bigl(s_{n_{0}},x_{0}\bigr)<\rho\bigl(x_{0},U_{i}^{c}\bigr)/2$. Thus, we have $x_{0}\in B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)\big/2\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)-\rho\bigl(x_{0},s_{n_{0}}\bigr)\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(s_{n_{0}},U_{i}^{c}\bigr)\bigr)=V_{n_{0},i}$. Therefore, we obtain that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$ for each $i\in I$.
@@
In order to prove Theorem 1.4, we need some notions. Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be an expanding Thurston map and $\mathcal{C}\subseteq\widehat{\mathbb{C}}$ be a Jordan curve containing $\operatorname{post}{f}$. For $n\in\mathbb{N}_{0}$, we define the set of $n$-tiles as
$\mathbf{X}^{n}(f,\mathcal{C})\coloneq\bigl\{\ \overline{A}\,:\,A\text{ is a connected component of }\widehat{\mathbb{C}}\smallsetminus f^{-n}(\mathcal{C})\bigr\}$
and define the set of $n$-edges as
$\mathbf{E}^{n}(f,\mathcal{C})\coloneq\bigl\{\ \overline{A}\,:\,A\text{ is a connected component of }f^{-n}(\mathcal{C})\smallsetminus f^{-n}(\operatorname{post}{f})\bigr\}.$
For $v\in\operatorname{crit}{f^{n}}$, let $W^{n}(v)$ denote the $n$-flower of $v$, as defined in [BM17, Definition 5.27]. By [BM17, Lemmas 5.28 (i) and 5.9 (v)], this set is explicitly given by
$W^{n}(v)=\mathopen{\{}v\mathclose{\}}\cup\bigcup\mathopen{\{}\operatorname{int}X:X\in\mathbf{X}^{n}(f,\mathcal{C}),\,v\in X\mathclose{\}}\cup\bigcup\mathopen{\{}\operatorname{int}e:e\in\mathbf{E}^{n}(f,\mathcal{C}),\,v\in e\mathclose{\}}.$
Note that flowers are open (in the standard topology on $\widehat{\mathbb{C}}$). Let $\overline{W}^{n}(v)$ be the closure of $W^{n}(v)$.
Remark 6.10.
For each $n\in\mathbb{N}_{0}$ and each $v\in\operatorname{crit}{f^{n}}$, we have $\overline{W}^{n}(v)=X_{1}\cup X_{2}\cup\cdots\cup X_{m}$, where $m\coloneq 2\deg_{f^{n}}(v)$, and $X_{1},\,X_{2},\,\dots,\,X_{m}$ are all the $n$-tiles that contain $v$ as a vertex (see [BM17, Lemma 5.28]). In particular, $W^{n}(v)$ is a simply connected open set. Moreover, if $f(v)=v$ and $v\in\operatorname{crit}{f^{n-1}}$, then it follows from [BM17, Proposition 5.16] that
$f(W^{n}(v))=W^{n-1}(v).$ (6.6)
@@
Now we show that there exists $k\in\mathbb{N}$ such that $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\mathrm{post}\bigl(f^{km}\bigr)$. Indeed, by $\operatorname{card}\mathopen{(}\mathrm{Per}(f)\cap\operatorname{post}{f}\mathclose{)}<+\infty$, there exists $k\in\mathbb{N}$ satisfying that $f^{k}(x)=x$ for each $x\in\mathrm{Per}(f)\cap\operatorname{post}{f}$. Thus, since $\mathrm{Per}(f^{n})\cap\mathrm{post}(f^{n})\subseteq\mathrm{Per}(f)\cap\operatorname{post}{f}$ for each $n\in\mathbb{N}$, we have $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\operatorname{post}{f^{km}}$.
Furthermore, by [BM17, Theorem 15.1] we can find $m\in\mathbb{N}$ such that there exists an $(f^{m})$-invariant Jordan curve $\mathcal{C}\subseteq S^{2}$ with $\mathrm{post}\bigl(f^{km}\bigr)=\operatorname{post}{f}\subseteq\mathcal{C}$. We set $F\coloneqq f^{km}$ and fix such a Jordan curve $\mathcal{C}$. Then by the previous discussion, we obtain that $F(\mathcal{C})\subseteq\mathcal{C}$ and all points in $\mathrm{Per}(F)\cap\operatorname{post}{F}$ are fixed points of $F$.

### v2

@@ 5.1. Jacobian and the transfer operator
Finally, we show that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. We argue this for $n$ by contradiction and consider a measure $\mu\in\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}$ with $\mu\neq\mu_{n}$. Then it is not hard to derive from (5.28) that $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$. Note that $\operatorname{card}\mathopen{(}C_{n}\smallsetminus Y_{n}\mathclose{)}<+\infty$ by item (iv) in the Assumptions. Thus, since $\mu\in\mathcal{P}(X;C_{n})\smallsetminus\mathcal{P}(X;C_{n}\cap Y_{n})$, we have $\mu(\{x_{0}\})>0$ for some $x_{0}\in C_{n}\smallsetminus Y_{n}$. Then we show that $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. By the assumptions in Theorem 5.10, $p_{n,i}\in U_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Hence, by $\mu\in\mathcal{K}_{n}$, it follows from (5.27) that $0\leqslant\mu(\{p_{n,i}\})\leqslant\mu(U_{n,i,k})\leqslant t_{n,i,k}$ for all $(n,i)\in H_{n}$ and $k\in\mathbb{N}$. Note that $\inf_{k\in\mathbb{N}}t_{n,i,k}=0$ for each $(n,i)\in H_{n}$ by (5.26). Then we have $\mu(\{p_{n,i}\})=0$ for each $(n,i)\in H_{n}$. Hence, by item (v) in the Assumptions, we obtain that $\mu(C_{n}\smallsetminus X_{n})=\mu(B_{n})=0$. Note that $\mu\in\mathcal{P}(X;C_{n})\cap\mathcal{M}(X,T_{n};X_{n})$ by $\mu\in\mathcal{K}_{n}$ and (5.27). Then by (1.1), we have $\mu\in\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n};X_{n})=\mathcal{P}(X;X_{n})\cap\mathcal{M}(X,T_{n})$. Hence, by Lemma 5.9, it follows from $\mu(\{x_{0}\})>0$ that $x_{0}$ is a periodic point of $T_{n}$ in $C_{n}\cap(X_{n}\smallsetminus Y_{n})$. Moreover, by $\mu\in\mathcal{P}(X;C_{n}\cap X_{n})$, the orbit of $x_{0}$ for $T_{n}$ is contained in $C_{n}\cap X_{n}$. Thus, $x_{0}\in B_{n}$ by the definition of $B_{n}$. This implies that $\mu(B_{n})\geqslant\mu(\{x_{0}\})>0$, which contradicts with $\mu(B_{n})=0$. Hence, we obtain that $\mathcal{M}(X,T_{n};Y_{n},J_{n})\cap\mathcal{K}_{n}\subseteq\{\mu_{n}\}$. This completes the proof of Claim 2.
@@
We now consider dynamical systems satisfying the hypotheses of Theorem 1.3 and the Assumptions whose maps $T_{n}$ are uniformly contracting near all periodic orbits in $B_{n}$ for $T_{n}$.
@@
    There exists a computable function $m\colon H\rightarrow\mathbb{N}$ and two sequences $\{r_{n,i}\}_{(n,i)\in H}$ and $\{\lambda_{n,i}\}_{(n,i)\in H}$ of uniformly computable real numbers such that for each $(n,i)\in H$, $m(n,i)$ is the period of $p_{n,i}$ for $T_{n}$, $r_{n,i}>0$, $0<\lambda_{n,i}<1$, and
    $\rho\bigl(p_{n,i},T_{n}^{m(n,i)}(q)\bigr)\leqslant\lambda_{n,i}\cdot\rho(p_{n,i},q)\quad\text{ for each }q\in B(p_{n,i},r_{n,i}).$
@@
We establish that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$ for each $(n,i)\in H$. We consider an arbitrary pair $(n,i)\in H$ and define $B(l)\coloneqq T_{n}^{lm(n,i)}(B(p_{n,i},r_{n,i}))$ for each $l\in\mathbb{N}_{0}$.
By item (iv) in the Assumptions, it follows from $C_{n}\subseteq X_{n}$ that $p_{n,i}\in B_{n}\subseteq C_{n}\cap(X_{n}\smallsetminus Y_{n})=C_{n}\smallsetminus Y_{n}$. Note that $\mu_{n}\in\mathcal{P}(X;C_{n}\cap Y_{n})$ by item (iv) in the Assumptions. Then we have $\mu_{n}(\{p_{n,i}\})=0$. Moreover, by the definition of $\{B(l)\}_{l\in\mathbb{N}_{0}}$, it follows from $\mu_{n}\in\mathcal{E}_{0}(T_{n},\phi_{n};Y_{n})\subseteq\mathcal{M}(X,T_{n})$ that $\mu_{n}(B(0))\leqslant\mu_{n}(B(l))$ for each $l\in\mathbb{N}_{0}$. Since $\lambda_{n,i}\in(0,1)$, we have $B(l)\subseteq B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\subseteq B(0)$ for each $l\in\mathbb{N}_{0}$. Thus, we have $\mu_{n}(B(0))=\mu_{n}\bigl(B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=\mu_{n}(B(l))$, hence, $\mu_{n}\bigl(B(0)\smallsetminus B\bigl(p_{n,i},r_{n,i}\lambda_{n,i}^{l}\bigr)\bigr)=0$. Letting $l\rightarrow+\infty$, we obtain that $\mu_{n}(B(0)\smallsetminus\{p_{n,i}\})=0$. Combined with $\mu_{n}(\{p_{n,i}\})=0$, this proves that $\mu_{n}(B(p_{n,i},r_{n,i}))=0$.
@@ 6. Computability of equilibrium states for expanding Thurston maps
In this section, we consider a class of nonuniformly expanding maps on the topological $2$-sphere known as expanding Thurston maps. There has been active research on both thermodynamic formalism for expanding Thurston maps (see e.g. [BM10, BM17, HP09, Li18, Li15, Li17, LS24b]) and algorithmic aspects of these maps (see e.g. [SY15, RSY20]).
We first review the definition of expanding Thurston maps, along with some key concepts and results. Then in Subsection 6.2, we focus on Misiurewicz–Thurston rational maps and apply Approach I (see Theorem 1.1) to establish Theorem 1.2. Finally, in Subsection 6.3, we study computable expanding Thurston maps whose critical points are computable. By Theorem 6.4, their measure-theoretic entropy functions may not be upper semicontinuous, which prevents us from using Approach I (see Theorem 1.1) to demonstrate the computability of their corresponding equilibrium states. Instead, we apply Approach II (see Theorem 1.3) to prove Theorem 1.4.
@@ 6.1. Expanding Thurston maps
$\lim_{n\to+\infty}\sup\bigl\{\operatorname{diam}_{d}X:X\text{ is a connected component of }f^{-n}\bigl(S^{2}\smallsetminus\mathcal{C}\bigr)\bigr\}=0.$
@@
The following result regarding the upper semicontinuity of the measure-theoretic entropy function is established in [LS24b, Theorem 1.1], extending [Li15, Corollary 1.3].
@@
Let $f\colon S^{2}\rightarrow S^{2}$ be an expanding Thurston map. Then the measure-theoretic entropy function of $f$ is upper semicontinuous if and only if $f$ has no periodic critical points.
@@ 6.2. Misiurewicz–Thurston rational maps
For all $n,\,m\in\mathbb{N}$, $\phi,\,u\in C\bigl(\widehat{\mathbb{C}},\sigma\bigr)$, a Misiurewicz–Thurston rational map $f=h_{1}/h_{2}$, and $x\in\widehat{\mathbb{C}}\smallsetminus\bigl\{f^{i}(\infty):1\leqslant i\leqslant m\bigr\}$, the algorithm outputs a rational $2^{-n}$-approximation of $\mathcal{L}^{m}_{\phi}(u)(x)$, given the following input data:
@@
By [Ly83, Corollary 1, p. 379], the measure-theoretic entropy function $\nu\mapsto h_{\nu}(f)$ is upper semicontinuous. By the hypothesis of Theorem 1.2, we have $\mathcal{E}(f,\phi_{n})=\{\mu_{n}\}$ for each $n\in\mathbb{N}$. Hence, to prove Theorem 1.2, by Theorem 1.1, it suffices to verify that the sequence $\{\phi_{n}\}_{n\in\mathbb{N}}$ of functions $\phi_{n}\colon\widehat{\mathbb{C}}\rightarrow\mathbb{R}$ satisfies properties (i) and (ii) in Theorem 1.1 in the case where $T_{n}\coloneqq f$ for each $n\in\mathbb{N}$.
@@
By (6.4), $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\lvert}q_{k,t}\mathclose{\rvert}$ for each $k\in\mathbb{N}$. Hence, by Definition 3.12, the function $F\colon\mathbb{N}\rightarrow\mathbb{R}$ given by $F(k)\coloneqq\sum_{t=1}^{l_{k}}2^{m_{k,t}-1}m_{k,t}\mathopen{\lvert}q_{k,t}\mathclose{\rvert}$ for each $k\in\mathbb{N}$ is a computable function satisfying $F(k)\geqslant\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}$ for each $k\in\mathbb{N}$. Thus, by (6.3), we have $\psi_{k}\in C^{0,\beta}\bigl(\widehat{\mathbb{C}},d_{v}\bigr)$ and $\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{\beta,d_{v}}\leqslant C\mathopen{\lvert}\psi_{k}\mathclose{\rvert}_{1,\sigma}\leqslant CF(k)$ for each $k\in\mathbb{N}$. Note that $C\in\mathbb{Q}$ and $F$ is a computable function. It follows from Proposition 6.8 that $\{P(f,\psi_{k})\}_{k\in\mathbb{N}}$ is a sequence of uniformly computable real numbers. This completes the proof of the claim.
@@ 6.3. Computable expanding Thurston maps
Before the proof of Theorem 1.4, we establish the following result.
@@
Let $(X,\,\rho,\,\mathcal{S})$ be a computable metric space in which $X$ is recursively compact. Suppose all balls in $(X,\rho)$ are connected. Assume that $I$ is a recursively enumerable set and that $\{U_{i}\}_{i\in I}$ is uniformly lower semi-computable open. Then there exists a recursively enumerable set $E\subseteq\mathbb{N}\times I$ and a sequence $\{V_{n,i}\}_{(n,i)\in E}$ of uniformly lower semi-computable open sets that are connected such that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$, where $E_{i}\coloneqq\{(n,i)\in E:n\in\mathbb{N}\}$ for each $i\in I$.
@@
Now we fix $i\in I$ and check that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$. Indeed, for each $(n,i)\in E_{i}$, since $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)$ is connected and $s_{n}\in U_{i}$, we have $B\bigl(s_{n},\rho\bigl(s_{n},U_{i}^{c}\bigr)\bigr)\subseteq U_{i}$. By the definition of $\{V_{n,i}\}_{(n,i)\in E}$, this implies that $\bigcup_{(n,i)\in E}V_{n,i}\subseteq U_{i}$. Then we establish $U_{i}\subseteq\bigcup_{(n,i)\in E_{i}}V_{n,i}$. We consider an arbitrary $x_{0}\in U_{i}$, and show that $x_{0}\in V_{n_{0},i}$ for some $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$. Since $\{s_{n}:n\in\mathbb{N}\}$ is dense in $X$ and $\{s_{n}:(n,i)\in E_{i}\}=\{s_{n}:n\in\mathbb{N}\}\cap U_{i}$, there exists $n_{0}\in\mathbb{N}$ with $(n_{0},i)\in E_{i}$ such that $\rho(s_{n_{0}},x_{0})<\rho\bigl(x_{0},U_{i}^{c}\bigr)/2$. Thus, we have $x_{0}\in B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)\big/2\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(x_{0},U_{i}^{c}\bigr)-\rho\bigl(x_{0},s_{n_{0}}\bigr)\bigr)\subseteq B\bigl(s_{n_{0}},\rho\bigl(s_{n_{0}},U_{i}^{c}\bigr)\bigr)=V_{n_{0},i}$. Therefore, we obtain that $U_{i}=\bigcup_{(n,i)\in E_{i}}V_{n,i}$ for each $i\in I$.
@@
To prove Theorem 1.4, we need to recall some notions. Let $f\colon\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ be an expanding Thurston map and $\mathcal{C}\subseteq\widehat{\mathbb{C}}$ be a Jordan curve containing $\operatorname{post}{f}$. For $n\in\mathbb{N}_{0}$, we define the set of $n$-tiles as
$\mathbf{X}^{n}(f,\mathcal{C})\coloneq\bigl\{\overline{A}:A\text{ is a connected component of }\widehat{\mathbb{C}}\smallsetminus f^{-n}(\mathcal{C})\bigr\}.$
For $n\in\mathbb{N}_{0}$ and $v\in f^{-n}(\operatorname{post}{f})$, we define the closed $n$-flower of $v$ as
$\overline{W}^{n}(v)\coloneqq\bigcup\mathopen{\{}X:X\in\mathbf{X}^{n}(f,\mathcal{C}),\,v\in X\mathclose{\}}.$
@@
Now we show that there exists $k\in\mathbb{N}$ such that $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\mathrm{post}\bigl(f^{km}\bigr)$. Indeed, since $\operatorname{card}\mathopen{(}\mathrm{Per}(f)\cap\operatorname{post}{f}\mathclose{)}<+\infty$, there exists $k\in\mathbb{N}$ such that for all $x\in\mathrm{Per}(f)\cap\operatorname{post}{f}$, we have $f^{k}(x)=x$. Thus, since $\mathrm{Per}(f^{n})\cap\mathrm{post}(f^{n})\subseteq\mathrm{Per}(f)\cap\operatorname{post}{f}$ for each $n\in\mathbb{N}$, we have $f^{km}(x)=x$ for all $m\in\mathbb{N}$ and $x\in\mathrm{Per}\bigl(f^{km}\bigr)\cap\operatorname{post}{f^{km}}$.
Furthermore, by [BM17, Theorem 15.1] we can find $m\in\mathbb{N}$ such that there exists an $f^{m}$-invariant Jordan curve $\mathcal{C}\subseteq S^{2}$ with $\mathrm{post}\bigl(f^{km}\bigr)=\operatorname{post}{f}\subseteq\mathcal{C}$. We set $F\coloneqq f^{km}$ and fix such a Jordan curve $\mathcal{C}$. Then by the previous discussion, we obtain that $F(\mathcal{C})\subseteq\mathcal{C}$ and all points in $\mathrm{Per}(F)\cap\operatorname{post}{F}$ are fixed points of $F$.